2. Convert the following hexadecimal numbers into decimal and binary (32 bit size): (a) A33 (b) ACDC
Answer #1

2) a) (A 33) Her oc). A- 1010 3 70110 100001010 0110 0110), io into decimal A 3 162 161 ( 3 160 lox16²+ 3x16 + 3 = 2611 ( 201(ЧЧ2- S-). - t) + 91W 81 +7 (37) +371 361 11 - 1 14 13 12 (o с 2 = . . tooth lo11ъотото) ( 1011) no? oole 0101rto *?(з дэ ө п